Iodine-131 is used in medical images for CT scans with contrast. Iodine-131 decay is first order...

  1. Iodine-131 is used in medical images for CT scans with contrast. Iodine-131 decay is first order with a half-life of 8.0 days. If a patient is administered a typical dose of 370 mg Iodine, how much Iodine-131 will remain after 24.0 days?
Answer #1

The answer to this question is quite simple

We don't even have to use any formula for this one

We have half life = 8 days

And we are checking it after 24 days

So we have 3 half lives in that period

One half life means the concentration becomes 1/2 of original concentration

Second half life = 1/4 of original concentration

And 3 rd half life = 1/8 of original concentration

So after 24 days remaining iodine 131 is

= 1/8 × 370 mg

= 46.25 gram

I hope this helps. If you have any query or want more detailed explanation, feel free to ask in the comments section below.